# Numerical simulation and analysis of mixing enhancement due to chaotic advection using an adaptive approach for approximating the dilution index

### v1

@@ 2.2 Random walk to model diffusion
Hamiltonian flows, as the ones defined above, only describe advective processes. To also simulate a diffusive process, we need to add some stochastic dynamics to the flow. Langevin and Einstein have shown how a random walk can be used to model the displacement of a Brownian particle Lemons and Gythiel (1997); Einstein (1906). For our application, we assume a mean square displacement of $\sigma^{2}t$ after time $t$ where $\sigma$ is the diffusion coefficient. Given the advective flow $\Psi^{\Delta t}$, the diffusion is added in terms of
@@
It is shown in Tomé and De Oliveira (2015) that the probability density of a particle following this modified flow does mirror the solution of the advection-diffusion equation with diffusion tensor $D=\frac{1}{2}\left(\begin{smallmatrix}\sigma^{2}&0\\ 0&\sigma^{2}\end{smallmatrix}\right)$.
@@ 2.3 Dilution index
Our final goal is to analyze how well the injected tracer dilutes within the system. Due to the stochastic nature of the flow $\Psi_{\sigma}$, defined in (11), the position of a particle after time $t>0$ is not deterministic. It is described by a probability density function $\rho:\Omega\rightarrow\mathbb{R}$, which can also be seen as the normalized concentration distribution of a plume. Distinguishing between spreading and dilution is essential in this context, as only the latter is associated with an increase in the volume of fluid occupied by the solute. An approach for measuring the dilution of the tracer is the dilution index introduced by Kitanidis Kitanidis (1994). It is defined as the exponential over the differential entropy of $\rho$ over $\Omega$
[…]
If the probability density function $\rho(x,t)$ represents a solution to the advection-diffusion equation with positive definite diffusion tensor, then $\frac{d\ln(\mathcal{E})}{dt}>0$ Kitanidis (1994).
@@
However, in our application, the exact representation of $\rho$ is not available. We compute an approximation by generating multiple realizations of the stochastic particle evolution defined in (11). We estimate the probability $P_{k}$ of the particle to be contained in cell $k$ as the number of realizations contained in cell $k$ divided by the total amount of realizations. It follows $\rho_{k}\approx\frac{P_{k}}{\Delta V_{k}}$. The approximated dilution index on a nonuniform grid is therefore given by
@@
with $H=0$ if and only if there exists one cell that contains all particles and $H=\ln(n)$ if and only if the particles are uniformly distributed over all cells Brémaud (2017). This directly implies
@@
for the approximated dilution index $E$. In the case of an unbounded domain, the upper limit cannot be reached, while for a bounded domain $\Omega$, the maximal dilution index $E_{\text{max}}=n\Delta V=\text{Area}(\Omega)$ is given as the size of the domain. For a bounded domain, it is convenient to normalize the dilution index by its maximum, yielding the reactor ratio Kitanidis (1994)
@@
However, using more grid cells than particles only makes sense if large parts of the domain are particle-free. It is important to notice that inequality (17) is only defined for the approximated dilution index $E$, which depends on the number of particles $m$ and a grid size $\Delta V$. Considering $\Delta V\rightarrow 0$ and keeping $m$ fixed, we yield $E\rightarrow 0$. On the contrary, the approximation of $\mathcal{E}$ in (13) using the exact values of $\rho_{k}$ will converge to $\mathcal{E}$ as $\Delta V\rightarrow 0$. Yielding $E\rightarrow\mathcal{E}$ is only possible for $\Delta V\rightarrow 0$ simultaneous to $m\rightarrow\infty$. Given that $m$ is fixed, selecting a suitable grid size can be compared to the problem of determining an appropriate size of the representative elementary volume (REV) to get the volumetric porosity of a porous medium Bear (1988). The volume of the REV shall be much smaller than the size of the domain to ensure a low approximation error in (13) and large enough to include a sufficient amount of particles, ensuring a low error in the approximation of $\rho_{k}$.
@@ 3.1 Numerical implementation of the pulsed source-sink system
As motivated in section 2.1.1, sink and source can be placed at ±$\pm\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)$ without loss of generality. However, the pulsed source-sink system is unbounded by construction, which is impractical for our application, as this would imply the necessity of an infinite amount of particles. We, therefore, define a bounded domain $[-2,2]^{2}$ with periodic-like boundary conditions. Similar to the standard periodic boundary condition, the periodic-like boundary condition can be motivated by a configuration where sources and sinks are repeatedly placed in the $\mathbb{R}^{2}$ space. Instead of simply repeating the configuration, source and sink are flipped above and below the domain. We show the resulting configuration of sources and sinks in Fig. 2(a). For a particle leaving the domain at the top or bottom, we specify
$\begin{split}x_{\text{new}}=-x,\qquad y_{\text{new}}=\begin{cases}y-4&y>0\\ y+4&y<0\end{cases}.\end{split}$
@@
$\begin{split}x_{\text{new}}=\begin{cases}x-4&x>0\\ x+4&x<0\end{cases},\qquad y_{\text{new}}=-y.\end{split}$
@@
Benchmarking the evaluation time of $\Psi_{\sigma}$ for the pulsed source-sink system when tracking a particle for time $\Delta t$ on an Intel i5 processor yields about 0.020.020.02\,0.02ms. As the evaluation of the advective particle motion in (3) is relatively cheap, we use $10$ million particles to achieve high accuracy results in our numerical experiments.
To parametrize the system, we have one parameter $\Lambda\in(0,\infty)$. Its definition in (2) requires $a$, $Q$ and $\tau$. Thereby, different combinations may result in the same behavior of the system. We hence fix $a=1$ and $Q=4\pi$ for all experiments. This yields $\Lambda=\sqrt{4\tau}$, which is modified by different choices of $\tau$. The choice of $Q$ is motivated by doubling the areal discharge implicitly used for source and sink in the RPM flow. Given $t_{\text{max}}=20$ in our experiments, we can determine the number of strokes of source and sink by $\frac{t_{\text{max}}}{2\tau}=\frac{10}{\tau}$. Given that all particles with distance $<\Lambda$ to the sink are extracted during the next stroke of the sink, we need to ensure that $\Lambda$ is sufficiently small such that the circles depicted in Fig. 1 have some distance to the domain boundary. We hence restrict $\Lambda^{2}$ to the interval $(0,0.5]$. The final selection for the experiment is $\Lambda^{2}\in\{0.1,0.2,0.3,0.4,0.5\}$. For completeness, Fig. 3 shows Poincaré sections for these parameters. Poincaré sections show the locations of particles tracked after each period of operation. In our case, these plots were produced by tracking five particles that started inside the chaotic region of the domain and saving their location after each time $2\tau$. Tracking a particle corresponds to repeatedly applying the Poincaré map (4). We observe particle-free island inside the plots, which are the KAM islands. These islands consist of periodic points with various periods Aref et al. (2017). They originate from the neighborhood of elliptic points and are invariant under the advective flow $\Psi$ Ottino (1989); Lester et al. (2009); Chaté et al. (1999). Due to the periodic nature of the islands, particles that start inside the chaotic region of the domain cannot enter these islands. Therefore, these islands stay free of particles in the Poincaré sections.
@@ 3.2 Numerical implementation of the RPM flow
In this setup, the boundary condition is less critical than in the previous case. Unlike the PPS system, where particles leave the domain due to the strokes of the source, particles in the RPM flow can leave the system only due to diffusion. Thus, introducing standard reflective boundary for our domain is sufficient. When computing the advective particle movement in the RPM flow, solving equation (8) is challenging. Since (6) cannot be solved for $\theta$ in terms of $(t_{\text{adv}},\psi)$, we perform a numerical inversion over a high-resolution grid, as proposed in Lester et al. (2009). The results in this paper use an equidistant grid with $000\times 1,$ 000 × 1 , 000 nodes in the $\{\theta,\psi\}$ space and a piecewise linear barycentric interpolation to approximate the function $t_{\text{adv}}^{-1}(t_{\text{adv}},\psi)$. To verify this approach, we computed the error approximation $|\theta-(t_{\text{adv}}^{-1}\circ t_{\text{adv}})(\theta,\psi)|$ using $10,000$ uniformly distributed random samples in the right half sphere, yielding a mean error in $\mathcal{O}(10^{-5})$. The same analysis was also performed with $100\times 100$ and $10\times 10$ nodes, yielding a mean error in $\mathcal{O}(10^{-3})$ and $\mathcal{O}(10^{-1})$, indicating that the approximation error decreases linearly with the number of nodes. In addition to an initialization time of about 303030\,30s to approximate $t_{\text{adv}}^{-1}$ on an Intel i5 processor, an evaluation of $\Psi_{\sigma}$ for the RPM flow takes about 0.20.20.2\,0.2ms, ten times slower than for the pulsed source-sink system. To balance computational accuracy and runtime, we decided to reduce the number of particles in our numerical experiments to one million. The results are less accurate than those of the pulsed source-sink system, but the numerical errors are sufficiently small to draw conclusions.
To control the system, we set the rotation angle $\Theta$ and the time $\tau$ between rotations. As shown in Lester et al. (2009), the parameter space is restricted to $\{\Theta,\tau\}\in[0,\pi]\times[0,\infty)$. Setting formally $\tau<0$ results in switching the position of the source and the sink, which is equivalent to a rotation of the domain by $180$ degree. In general, we can restrict $\Theta\in[-\pi,\pi]$. Additionally, the reflection symmetry of the base flow shows the equivalence of $-\Theta$ to $\Theta$, which gives the desired result. Alike for the pulsed source-sink system, we use $t_{\text{max}}=20$, restricting the $\tau$ parameter space further to $(0,t_{\text{max}}]$. For the experiments, we use the Cartesian product over the two sets of parameters $\Theta\in\{0,\frac{\pi}{6},\frac{\pi}{3},\frac{\pi}{2},\frac{2\pi}{3},\frac{5\pi}{6}\}$ and $\tau\in\{0.2,0.5\}$. The Poincaré sections in Fig. 4 were created by repeatedly applying the Poincare map (10) on three particles that started in the chaotic region. We again see KAM-islands filling big parts of the domain. Compared to the pulsed source-sink system, the percentage of the domain covered by islands is bigger, while the absolute size might be equal or even smaller due to the difference in the domain size. We also observe more differences between the results for each parameter configuration. For $\Theta=0$, the source and sink do not rotate, and the particles always stay on the trajectory in the reference system they are initially placed on. In this configuration, no chaotic flow structures can be observed. For $\Theta>0$, we observe KAM islands. The size of the region covered by islands differs for each parameter configuration. As we observe chaotic advection only between the KAM islands in the chaotic region, the KAM island structure particularly impacts the chaotic mixing of the flow. We expect increased mixing for $\Theta=\frac{1}{6}\pi$ and $\tau=0.5$, where we only find three very small islands. On the contrary, for $\Theta=\frac{1}{6}\pi$ and $\tau=0.2$, nearly half of the domain is particle-free in Fig. 4, indicating less mixing enhancement by chaotic advection.
@@
Figure 4: Poincaré sections of the RPM flow for $\Theta\in\left\{0,\frac{1}{6}\pi,\frac{1}{3}\pi,\frac{1}{2}\pi,\frac{2}{3}\pi,\frac{5}{6}\pi\right\}$ and $\tau\in\{0.2,0.5\}$. White areas not occupied by particles are the KAM islands.
@@ 3.4 Diffusion coefficient
In order to compare the PPS and the RPM systems, we define a relationship between the diffusion coefficient $\sigma$ based on the Péclet number. Therefore, the results of the systems are comparable, independent of the different domain sizes and flow velocities. The Péclet number is defined as $\text{Pe}=\frac{Lu}{\sigma^{2}}$, with characteristic length $L$, flow velocity $u$ and diffusion coefficient $\sigma$. For the pulsed source-sink system, we alternate between $v_{-}$ for operating the sink and $v_{+}$ for operating the source. Using $L_{\text{PSS}}=4$ and
$\begin{split}u_{\text{PSS}}&=\int_{[-2,2]^{2}}\left|\left|v_{-}\right|\right|\,d(x,y)\\ &=\int_{[-2,2]^{2}}\left|\left|v_{+}\right|\right|\,d(x,y)\approx\frac{Q}{2\pi}13.342\end{split}$
@@
$u_{\text{RPM}}=\int_{B_{1}\left(\begin{smallmatrix}0\\ 0\end{smallmatrix}\right)}\left|\left|v_{\text{RPM}}\right|\right|\,d(x,y)\approx 7.328$
and $L_{\text{RPM}}=\sqrt{\pi}$ yield $\text{Pe}_{\text{RPM}}\approx\frac{12.989}{\sigma^{2}}$. To match the Péclet numbers of both systems, we define a relationship between the diffusion coefficient $\sigma$, yielding
@@ 3.5 A computationally reliable approach for the approximation of the dilution index
The definition of the dilution index uses a zero-order estimation method to approximate the particle density in the domain, called nearest grid point method Peterka et al. (2016). Decreasing the grid size leads to approximation errors in the estimation of the mean probability density $\rho_{k}\approx\frac{P_{k}}{\Delta V_{k}}$ inside each cell $k=1,...,n$. For a given number of particles $m$ and an equidistant grid of grid size $\Delta V$, we have derived an upper bound for the dilution index $E$ in (17). Decreasing $\Delta V\rightarrow 0$ results in $E\leq m\Delta V\rightarrow 0$ independent of the underlying particle distribution. An unrealistic result. Given the distribution of $m$ particles, our goal is to establish a sophisticated approach to select a grid size sufficiently small to reflect fine structures in the plume but big enough to ensure a low error in the approximation of $\rho_{k}$. The obtained density approximation is called "fine-grained density" Krasnopolskaya and Meleshko (2009).
@@
We assume an instantaneous pulsed injection. To select an appropriate grid size, we compare the numerical approximation of the dilution index against its analytical solution. Given the diffusion coefficient $\sigma$, the plume is Gaussian distributed with variance $\sigma^{2}t$ at time $t$ Kitanidis (1994). We can compute the analytical dilution index by inserting the corresponding Gaussian density function $\rho(x,t)$ into (12), yielding
[…]
$\Delta_{\delta}[E](h):=\frac{\ln(E(e^{\delta}h))-\ln(E(h))}{\delta}.$ (19)
@@
Multiplying $h$ by $e^{\delta}$ is equivalent to taking a $\delta$ step from $h$, respecting the logarithmic scale. This approach has the advantage that $\Delta_{\delta}[E](h)\rightarrow 2$ as $h\rightarrow 0$. Comparing the derivative in Fig. 6(c) to the relative approximation error in Fig. 6(b), we see that the grid sizes of low approximation errors indeed lie at the center of the plateau. This observation motivates the choice of the grid size $h$ to be the minimum of the derivative. Independent of the used diffusion coefficient $\sigma$, we observe similar relative errors close to the optimal grid size $h^{*}$. The more particles used, the clearer the minimum.
@@
This approach also equations with the theory of REV used to compute a homogenization of the volumetric porosity of a porous medium. Lowering the size of the REV below a particular value $\Delta V_{0}$ results in large fluctuations in the ratio of void space in the representative volume compared to its size. As the size goes to zero, the ratio takes a value in $\{0,1\}$, depending on the centroid of the REV. Equal to 0 if the centroid is inside the pore, or equal to 1 if it is inside the solid matrix. This leads to an undesirable result of the homogenization. To determine the volumetric porosity, the limit $\Delta V_{i}\rightarrow\Delta V_{0}$ is used Bear (1988).
@@
The example of the pulse injection also demonstrates the importance of a sophisticated way of choosing a suitable grid size. Fixing one grid size for all diffusion coefficients leads to huge errors in some dilution index approximations. We also find the number of particles $m$ to be crucial. Using more particles increases the accuracy of the approximation of $\rho_{k}$. This allows for a smaller grid size $h^{*}$, resulting in a higher accuracy in the approximation.
[…]

### v2

@@ 2.2 Random walk to model diffusion
Hamiltonian flows, as the ones defined above, only describe advective processes. To also simulate a diffusive process, we need to add some stochastic dynamics to the flow. Langevin and Einstein have shown how a random walk can be used to model the displacement of a Brownian particle [25, 13]. For our application, we assume a mean square displacement of $\sigma^{2}t$ after time $t$ where $\sigma$ is the diffusion coefficient. Given the advective flow $\Psi^{\Delta t}$, the diffusion is added in terms of
@@
It is shown in [52] that the probability density of a particle following this modified flow does mirror the solution of the advection-diffusion equation with diffusion tensor $D=1/2\left(\begin{smallmatrix}\sigma^{2}&0\\ 0&\sigma^{2}\end{smallmatrix}\right)$.
@@ 2.3 Dilution index
Our final goal is to analyze how well the injected tracer mixes within the system. Due to the stochastic nature of the flow $\Psi_{\sigma}$, defined in (11), the position of a particle after time $t>0$ is not deterministic. It is described by a probability density function $\rho:\Omega\rightarrow\mathbb{R}$, which can also be seen as the normalized concentration distribution of a plume. Distinguishing between spreading and dilution is essential in this context, as only the latter is associated with an increase in the volume of fluid occupied by the solute. An approach for measuring the dilution of the tracer is the dilution index introduced by Kitanidis [20]. It is defined as the exponential over the differential entropy of $\rho$ over $\Omega$
$\mathcal{E}=\exp\left(-\int_{\Omega}\rho(x)\ln(\rho(x))\,dx\right)$ (12)
If the probability density function $\rho(x,t)$ represents a solution to the advection-diffusion equation with positive definite diffusion tensor, then $\frac{d\ln(\mathcal{E})}{dt}>0$ [20].
@@
However, in our application, the exact representation of $\rho$ is not available. We compute an approximation by generating multiple realizations of the stochastic particle evolution defined in (11). We estimate the probability $P_{k}$ of the particle to be contained in cell $k$ as the number of realizations contained in cell $k$ divided by the total amount of realizations. It follows $\rho_{k}\approx P_{k}/\Delta V_{k}$. The approximated dilution index on a nonuniform grid is therefore given by
@@
with $H=0$ if and only if there exists one cell that contains all particles and $H=\ln(n)$ if and only if the particles are uniformly distributed over all cells [9]. This directly implies
@@
for the approximated dilution index $E$. In the case of an unbounded domain, the upper limit cannot be reached, while for a bounded domain $\Omega$, the maximal dilution index $E_{\text{max}}=n\Delta V=\text{Area}(\Omega)$ is given as the size of the domain. For a bounded domain, it is convenient to normalize the dilution index by its maximum, yielding the reactor ratio [20]
@@
However, using more grid cells than particles only makes sense if large parts of the domain are particle-free. It is important to notice that inequality (17) is only defined for the approximated dilution index $E$, which depends on the number of particles $m$ and a grid size $\Delta V$. Considering $\Delta V\rightarrow 0$ and keeping $m$ fixed, we yield $E\rightarrow 0$. On the contrary, the approximation of $\mathcal{E}$ in (13) using the exact values of $\rho_{k}$ will converge to $\mathcal{E}$ as $\Delta V\rightarrow 0$. Yielding $E\rightarrow\mathcal{E}$ is only possible for $\Delta V\rightarrow 0$ simultaneous to $m\rightarrow\infty$. Given that $m$ is fixed, selecting a suitable grid size can be compared to the problem of determining an appropriate size of the representative elementary volume (REV) to get the volumetric porosity of a porous medium [7]. The volume of the REV shall be much smaller than the size of the domain to ensure a low approximation error in (13) and large enough to include a sufficient amount of particles, ensuring a low error in the approximation of $\rho_{k}$.
@@ 3.1 Numerical implementation of the pulsed source-sink system
As motivated in Section 2.1.1, sink and source can be placed at $\pm\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)$ without loss of generality. However, the pulsed source-sink system is unbounded by construction, which is impractical for our application, as this would imply the necessity of an infinite amount of particles. We, therefore, define a bounded domain $2,2]^{2}$ with periodic-like boundary conditions. Similar to the standard periodic boundary condition, the periodic-like boundary condition can be motivated by a configuration where sources and sinks are repeatedly placed in the $\mathbb{R}^{2}$ space. Instead of simply repeating the configuration, source and sink are flipped above and below the domain. We show the resulting configuration of sources and sinks in Fig. 2(a). For a particle leaving the domain at the top or bottom, we specify
$x_{\text{new}}=-x,\qquad y_{\text{new}}=\begin{cases}y-4&y>0\\ y+4&y<0\end{cases}.$
@@
$x_{\text{new}}=\begin{cases}x-4&x>0\\ x+4&x<0\end{cases},\qquad y_{\text{new}}=-y.$
@@
Benchmarking the evaluation time of $\Psi_{\sigma}$ for the pulsed source-sink system when tracking a particle for time $\Delta t$ on an Intel i5 processor yields about 0.020.02\,ms. This runtime restricts the number of particles used in the particle tracking experiments. Choosing a sufficient amount of particles is crucial to achieving a high resolution of the particle density approximation [41], especially for mixing processes limited by diffusion [37], as in our case. The number of particles $m$ limits the resolution of low-concentration regions as a grid cell can only contain an integer number of particles. Hence, the particle density of a non-empty cell $V_{k}$ can not be smaller than $\frac{1}{m\,\Delta V_{k}}$. Following the argumentation of [37], the simulation of each solute molecule in the system is impossible due to their enormous number. Therefore, we restrict ourselves to $m=10^{7}$ particles, which is inside their proposed range of $10^{6}$$10^{9}$.
To parametrize the system, we have one parameter $\Lambda\in(0,\infty)$. Its definition in (2) requires $a$, $Q$ and $\tau$. Thereby, different combinations may result in the same behavior of the system. We hence fix $a=1$ and $Q=4\pi$ for all experiments. This yields $\Lambda=\sqrt{4\tau}$, which is modified by different choices of $\tau$. The choice of $Q$ is motivated by doubling the areal discharge implicitly used for source and sink in the RPM flow. Given $t_{\text{max}}=20$ in our experiments, we can determine the number of strokes of source and sink by $t_{\text{max}}/(2\tau)=10/\tau$. Given that all particles with distance $\Lambda$ to the sink are extracted during the next stroke of the sink, we need to ensure that $\Lambda$ is sufficiently small such that the circles depicted in Fig. 1 have some distance to the domain boundary. We hence restrict $\Lambda^{2}$ to the interval $(0,0.5]$. The final selection for the experiment is $\Lambda^{2}\in\{0.1,0.2,0.3,0.4,0.5\}$. For completeness, Fig. 3 shows Poincaré sections for these parameters. Poincaré sections show the locations of particles tracked after each period of operation. In our case, these plots were produced by tracking five particles that started inside the chaotic region of the domain and saving their location after each time $2\tau$. Tracking a particle corresponds to repeatedly applying the Poincaré map (4). We observe particle-free island inside the plots, which are the KAM islands. These islands consist of periodic points with various periods [1]. They originate from the neighborhood of elliptic points and are invariant under the advective flow $\Psi$ [33, 26, 10]. Due to the periodic nature of the islands, particles that start inside the chaotic region of the domain cannot enter these islands. Therefore, these islands stay free of particles in the Poincaré sections.
@@ 3.2 Numerical implementation of the RPM flow
Figure 4: Poincaré sections of the RPM flow for $\Theta\in\left\{0,\pi/6,\pi/3,\pi/2,2\pi/3,5\pi/6\right\}$ and $\tau\in\{0.2,0.5\}$. White areas not occupied by particles are the KAM islands.
In this setup, the boundary condition is less critical than in the previous case. Unlike the PPS system, where particles leave the domain due to the strokes of the source, particles in the RPM flow can leave the system only due to diffusion. Thus, introducing standard reflective boundary for our domain is sufficient. When computing the advective particle movement in the RPM flow, solving equation (8) is challenging. Since (6) cannot be solved for $\theta$ in terms of $(t_{\text{adv}},\psi)$, we perform a numerical inversion over a high-resolution grid, as proposed in [26]. The results in this paper use an equidistant grid with $1,000\times 1,000$ nodes in the $\{\theta,\psi\}$ space and a piecewise linear barycentric interpolation to approximate the function $t_{\text{adv}}^{-1}(t_{\text{adv}},\psi)$. To verify this approach, we computed the error approximation $|\theta-(t_{\text{adv}}^{-1}\circ t_{\text{adv}})(\theta,\psi)|$ using $10,000$ uniformly distributed random samples in the right half sphere, yielding a mean error in $\mathcal{O}(10^{-5})$. The same analysis was also performed with $100\times 100$ and $10\times 10$ nodes, yielding a mean error in $\mathcal{O}(10^{-3})$ and $\mathcal{O}(10^{-1})$, indicating that the approximation error decreases linearly with the number of nodes. In addition to an initialization time of about 3030\,s to approximate $t_{\text{adv}}^{-1}$ on an Intel i5 processor, an evaluation of $\Psi_{\sigma}$ for the RPM flow takes about 0.20.2\,ms, ten times slower than for the pulsed source-sink system. To balance computational accuracy and runtime, we decided to reduce the number of particles in our numerical experiments to one million. The results are less accurate than those of the pulsed source-sink system, but the numerical errors are sufficiently small to draw conclusions.
To control the system, we set the rotation angle $\Theta$ and the time $\tau$ between rotations. As shown in [26], the parameter space is restricted to $\{\Theta,\tau\}\in[0,\pi]\times[0,\infty)$. Setting formally $\tau<0$ results in switching the position of the source and the sink, which is equivalent to a rotation of the domain by $180$ degree. In general, we can restrict $\Theta\in[-\pi,\pi]$. Additionally, the reflection symmetry of the base flow shows the equivalence of $\Theta$ to $\Theta$, which gives the desired result. Alike for the pulsed source-sink system, we use $t_{\text{max}}=20$, restricting the $\tau$ parameter space further to $(0,t_{\text{max}}]$. For the experiments, we use the Cartesian product over the two sets of parameters $\Theta\in\left\{0,\pi/6,\pi/3,\pi/2,2\pi/3,5\pi/6\right\}$ and $\tau\in\{0.2,0.5\}$. The Poincaré sections in Fig. 4 were created by repeatedly applying the Poincare map (10) on three particles that started in the chaotic region. We again see KAM-islands filling large parts of the domain. Compared to the pulsed source-sink system, the percentage of the domain covered by islands is bigger, while the absolute size might be equal or even smaller due to the difference in the domain size. We also observe more differences between the results for each parameter configuration. For $\Theta=0$, the source and sink do not rotate, and the particles always stay on the trajectory in the reference system they are initially placed on. In this configuration, no chaotic flow structures can be observed. For $\Theta>0$, we observe KAM islands. The size of the region covered by islands differs for each parameter configuration. As we observe chaotic advection only between the KAM islands in the chaotic region, the KAM island structure particularly impacts the chaotic mixing of the flow. We expect increased mixing for $\Theta=\pi/6$ and $\tau=0.5$, where we only find three very small islands. On the contrary, for $\Theta=\pi/6$ and $\tau=0.2$, nearly half of the domain is particle-free in Fig. 4, indicating less mixing enhancement by chaotic advection.
@@ 3.4 Diffusion coefficient
In order to compare the PPS and the RPM systems, we define a relationship between the diffusion coefficient $\sigma$ based on the Péclet number. Therefore, the results of the systems are comparable, independent of the different domain sizes and flow velocities. The Péclet number is defined as $\text{Pe}=Lu/\sigma^{2}$, with characteristic length $L$, flow velocity $u$ and diffusion coefficient $\sigma$. For the pulsed source-sink system, we alternate between $v_{-}$ for operating the sink and $v_{+}$ for operating the source. Using $L_{\text{PSS}}=4$ and
[…]
@@
[…]
and $L_{\text{RPM}}=\sqrt{\pi}$ yield $\text{Pe}_{\text{RPM}}\approx 12.989/\sigma^{2}$. To match the Péclet numbers of both systems, we define a relationship between the diffusion coefficient $\sigma$, yielding
@@ 3.5 A computationally reliable approach for the approximation of the dilution index
The definition of the dilution index uses a zero-order estimation method to approximate the particle density in the domain, called nearest grid point method [36]. Decreasing the grid size leads to approximation errors in the estimation of the mean probability density $\rho_{k}\approx\frac{P_{k}}{\Delta V_{k}}$ inside each cell k=1,,nk=1,...,n. For a given number of particles $m$ and an equidistant grid of grid size $\Delta V$, we have derived an upper bound for the dilution index $E$ in (17). Decreasing $\Delta V\rightarrow 0$ results in $E\leq m\Delta V\rightarrow 0$ independent of the underlying particle distribution. An unrealistic result. Given the distribution of $m$ particles, our goal is to establish a sophisticated approach to select a grid size sufficiently small to reflect fine structures in the plume but big enough to ensure a low error in the approximation of $\rho_{k}$. The obtained density approximation is called ”fine-grained density” [21].
@@
We assume an instantaneous pulsed injection. To select an appropriate grid size, we compare the numerical approximation of the dilution index against its analytical solution. Given the diffusion coefficient $\sigma$, the plume is Gaussian distributed with variance $\sigma^{2}t$ at time $t$ [20]. We can compute the analytical dilution index by inserting the corresponding Gaussian density function $\rho(x,t)$ into (12), yielding
$\begin{split}\mathcal{E}(t)=\exp\left(-\int_{\Omega}\rho(x,t)\ln(\rho(x,t))\,dx\right)=\exp(\ln(2\pi\sigma^{2}t)+1)=2\pi\sigma^{2}t\exp(1).\end{split}$ (18)
Figure S1 shows the results for $t=1$ for different diffusion coefficients and m=10 000m=10\,000 particles. When $\sigma=0$ (black color), one cell contains all particles and $E=\Delta V=h^{2}$; thus $E$ decreases quadratically with $h$ (Fig. S1(a)). As shown in (16), this is a general lower bound for $E$. When $\sigma>0$, for big grid sizes, the dilution index converges to the analytical dilution index (Fig. S1(b)). By decreasing the grid size, we reach an optimal grid size $h^{*}$ with error zero. Notice that Fig. S1 was created using only a finite selection of grid sizes that do not include $h^{*}$. Thus, the approximation error is always greater than zero. For smaller grid sizes, the error increases again due to increasing errors in the approximations of $\rho_{k}$, and the dilution index approaches the upper bound $m\Delta V$ as $h$ tends to 0 (Fig. S1(a)). This results in a plateau in the plot $E$ over $h$ at the value of the analytical dilution index. The length of the plateau varies depending on the number of particles used in the experiments. In general, at the center of the plateau, it holds that $\frac{dE}{dh}(h^{*})=0$. For a discrete selection of grid sizes, the one closest to $h^{*}$ may be found by estimating the minimum of the forward finite differences. This approach has the problem that with $h\rightarrow 0$, we yield $E\rightarrow 0$ quadratically and hence also $\frac{dE}{dh}(h)\rightarrow 0$. Hence, finding the minimal derivative of $E$ using forward finite differences over a discrete grid may select erroneously small grid sizes. Instead, we use the derivative of $\ln(E)$ with respect to the logarithmic scaled grid size $h$. This alternative term is also zero if and only if $h=h^{*}$. We compute the derivative using forward finite differences with step size $\delta>0$
$\Delta_{\delta}[E](h):=\frac{\ln(E(h\exp(\delta)))-\ln(E(h))}{\delta}.$ (19)
@@
Multiplying $h$ by $\exp(\delta)$ is equivalent to taking a $\delta$ step from $h$, respecting the logarithmic scale. This approach has the advantage that $\Delta_{\delta}[E](h)\rightarrow 2$ as $h\rightarrow 0$. Comparing the derivative in Fig. S1(c) to the relative approximation error in Fig. S1(b), we see that the grid sizes of low approximation errors indeed lie at the center of the plateau. This observation motivates the choice of the grid size $h$ to be the minimum of the derivative. Independent of the used diffusion coefficient $\sigma$, we observe similar relative errors close to the optimal grid size $h^{*}$. The more particles used, the clearer the minimum.
@@
This approach also aligns with the theory of REV used to compute a homogenization of the volumetric porosity of a porous medium. Lowering the size of the REV below a particular value $\Delta V_{0}$ results in large fluctuations in the ratio of void space in the representative volume compared to its size. As the size goes to zero, the ratio takes a value in $\{0,1\}$, depending on the centroid of the REV. Equal to 0 if the centroid is inside the pore, or equal to 1 if it is inside the solid matrix. This leads to an undesirable result of the homogenization. To determine the volumetric porosity, the limit $\Delta V_{i}\rightarrow\Delta V_{0}$ is used [7].
Figure 6: (a) Shows the value of the dilution index $E$ over the grid size $h$ for m=10 000m=10\,000 particles and different diffusion coefficients $\sigma$. The dotted line gives the upper bound of the dilution index $m\Delta V$. The continuous lightened lines help to visualize the plateaus forming at the value of the analytical dilution index. (b) Shows the relative error $\frac{|\mathcal{E}-E|}{\mathcal{E}}$ of the numerical dilution index $E$ depending on the grid size $h$ to the analytical value $\mathcal{E}$ from (18). We observe a convergence $E\rightarrow\mathcal{E}$, until we pass the optimal grid size $h^{*}$. For smaller $h$, the dilution index dissipates from the analytical value. (c) Shows the approximated derivative of the dilution index $\Delta_{\delta}[E]$, using forward finite differences in the log-log scale as defined in (19), for $\sigma\neq 0$. The minima of the derivatives correspond approximately to the minima in the approximation error of plot (b).
@@
The example of the pulse injection also demonstrates the importance of a sophisticated way of choosing a suitable grid size. Fixing one grid size for all diffusion coefficients leads to huge errors in some dilution index approximations. We also find the number of particles $m$ to be crucial. Using more particles increases the accuracy of the approximation of $\rho_{k}$. This allows for a smaller grid size $h^{*}$, resulting in a higher accuracy in the approximation. A more detailed figure that further supports our findings concerning the number of particles $m$ is included in the supplementary material section S2 Fig. S1.